### v2

# UniPart: Part-Level 3D Generation with Unified 3D Geom–Seg Latents

Abstract
Part-level 3D generation is essential for applications requiring decomposable and structured 3D synthesis. However, existing methods either rely on implicit part segmentation with limited granularity control or depend on strong external segmenters trained on large annotated datasets. In this work, we observe that part awareness emerges naturally during whole-object geometry learning and propose Geom-Seg VecSet, a unified geometry–segmentation latent representation that jointly encodes object geometry and part-level structure. Building on this representation, we introduce UniPart, a two-stage latent diffusion framework for image-guided part-level 3D generation. The first stage performs joint geometry generation and latent part segmentation, while the second stage conditions part-level diffusion on both whole-object and part-specific latents. A dual-space generation scheme further enhances geometric fidelity by predicting part latents in both global and canonical spaces. Extensive experiments demonstrate that UniPart achieves superior segmentation controllability and part-level geometric quality compared with existing approaches.
1 Introduction
With the rapid development of 3D content creation in fields such as computer graphics, robotics, virtual reality, and digital manufacturing, the demand for controllable and semantically structured 3D shape generation has increased significantly. Traditional 3D generation methods [42, 44, 16, 36] primarily focus on producing holistic meshes that capture only the overall geometry of objects. However, real-world applications—such as part-aware editing, physical simulation, robotic manipulation, and modular design—require not only accurate geometry but also a decomposable understanding of object parts and their relationships. This need has motivated increasing interest in part-level 3D shape generation, which aims to synthesize objects as sets of coherent and semantically meaningful components. Compared to holistic 3D reconstruction, part-level 3D shape generation poses two principal challenges. First, achieving structurally and semantically coherent part segmentation is inherently challenging. The model must infer an appropriate segmentation granularity while preserving both global structural integrity and semantic consistency. The wide variability in object geometries and part compositions across—and even within—categories further makes coherent and meaningful decomposition highly nontrivial. Second, part geometry degradation frequently occurs during generation, particularly for small or thin components. These fine structures tend to lose geometric fidelity due to limited generative resolution and the difficulty of jointly preserving global coherence and local geometric details.
Existing solutions develop implicit segmentation by composing groups of latent vectors for decomposable part generation [18, 31, 4], but suffer from limited controllability over segmentation granularity and suboptimal part-level geometric fidelity. Another line of work first performs whole-object generation and subsequently applies explicit segmentation using multi-view images [3], dense points [37], low-resolution voxels [40], or mesh faces [38]. These approaches achieve higher segmentation quality and controllability but require a powerful auxiliary segmenter trained on abundant labeled data. We begin with the observation that part awareness emerges implicitly during whole-geometry learning. This arises from two factors: (i) intra- and inter-part geometric correlations are implicitly promoted during VAE latent-space construction for efficient geometric compression, leading to part-aware latent representations; (ii) image conditions encoded by pretrained vision models such as DINOv2 [24] carry rich part-level semantic cues, which facilitate part-aware learning in image-to-geometry diffusion. Motivated by these observations, we propose to construct a unified 3D geometry–segmentation latent space based on the VecSet representation [42]. This latent space is learned to jointly encode geometric shape and part-segmentation information into Geom-Seg VecSet, and to decode each latent vector into its corresponding geometric contribution and part label. Thanks to the tight integration of part understanding and geometry learning, our representation maintains reconstruction quality comparable to the original VecSet.
For image-guided part-level 3D generation, we introduce a two-stage latent diffusion model, UniPart. We first perform whole-object diffusion to generate Geom-Seg VecSet, yielding both the global geometric latent and the corresponding part latent mask. Compared with methods that conduct part-level generation with implicit segmentation, our approach produces explicit part-latent masks, providing more flexible segmentation granularity and higher part-level controllability. Unlike previous generate-then-segment pipelines, our approach performs geometry generation and part segmentation jointly. This enables more efficient part learning by leveraging the rich part-aware priors implicitly captured in the geometry generator. Moreover, latent-space segmentation actually strikes a balance between granularity—akin to dense point-based approaches—and robustness—similar to low-resolution voxel methods. In the second stage, we directly use the segmented part latents, along with the global latent, as dual conditioning signals for part-level generation. Compared with prior approaches that require whole-object geometry generation and segmentation on explicit representations to condition the part-level diffusion [37, 38, 40], our latent-based conditioning offers three key advantages: (i) avoiding geometry degradation introduced by decode–encode cycles (e.g., decoding to SDF for lossy marching-cube mesh extraction and re-encoding sampled surface points [37, 38]); (ii) providing domain-aligned conditioning signals that facilitate latent-space diffusion learning; (iii) injecting informative part–whole relational cues inherently preserved in the latent space. Furthermore, our part-level diffusion performs a novel dual-space latent generation procedure that jointly predicts part geometry latents in both the global coordinate space and a normalized canonical space. Decoding these two latents provides the part’s canonical geometry (from the normalized space latent) and its scale and placement within the object (from the global-space latent), which substantially mitigates part-geometry degradation and enables seamless and consistent part composition. Extensive experiments are conducted to validate the effectiveness and superiority of our UniPart on both part segmentation and geometry quality.
Contributions
Our key contributions are as follows:
  • We introduce Geom-Seg VecSet, a unified geometry-segmentation latent representation that enables efficient part-level understanding directly from whole-object generative learning.
  • We develop UniPart, a two-stage latent diffusion model that first performs geometry generation and part-latent segmentation, and then conducts dual-space part diffusion for seamless part composition and full-resolution part geometry, conditioned on both part and whole latents.
  • We conduct extensive experiments demonstrating the superiority of our method for part-level 3D object generation.
2 Related Work
Object-level 3D Generation
Recent object-level 3D generation methods fall into two main categories: 2D-lifting and native 3D approaches. DreamFusion [26] pioneered the use of 2D diffusion priors via Score Distillation Sampling [1, 33] to optimize 3D representations, while follow-ups [22, 34, 21, 28] improve multi-view consistency by generating aligned image sets before reconstruction. In parallel, native 3D generative models are trained directly on large-scale datasets [10, 8]. CLAY [42] and related works [44, 16, 15] adopt a VecSet-based VAE–LDM framework [41] to generate meshes and require high-quality watertight meshes for training. Trellis [35] introduces an explicit voxel-based latent space trained with rendering losses, offering stronger geometric inductive bias [17, 13]. However, scaling such representations to high resolution remains a key challenge for fine-detail synthesis. Our work operates at the part level and builds upon the VecSet paradigm: we retain the VecSet-based latent architecture but extend the purely geometric latent space by incorporating geometry-aligned segmentation features, thereby constructing a novel unified geometry–segmentation latent space that enables structured part decomposition while preserving reconstruction fidelity.
3D Part Segmentation
Recent part segmentation methods leverage either 2D priors or native 3D supervision. Approaches such as SaMesh [30] lift 2D SAM [14, 27] predictions to 3D via multi-view post-processing, while SAMPart3D [39] and PartField [20] distill SAM priors into a 3D semantic feature field for clustering-based decomposition. In contrast, 3D native methods like P3-SAM [23] and PartSAM [45] are trained directly on large-scale part-annotated meshes and produce point-wise semantic features for mask prompting. GeoSAM2 [11] adopts a hybrid strategy: it finetunes the pretrained SAM on 3D data to enable multi-view consistent segmentation, then back-projects the 2D segmentation onto the mesh surface to obtain a coherent 3D part decomposition. Unlike these strategies that conduct segmentation on explicit 2D/3D representations, our method performs decomposition on the latent representation. By integrating high-dimensional, geometry-aligned semantic features into the latent representation, our approach inherits the robustness and boundary-awareness of sampling while enabling end-to-end part-level generation without relying on hard segmentation masks. Meanwhile, solving the segmentation task as a diffusion-sampling task is much more reasonable than regression. Unlike regression, which forces a single deterministic output, treating part segmentation as a diffusion-based sampling task better captures the inherent ambiguity and multi-modality of valid part decompositions.
Part-level 3D Generation
Existing part-level 3D generation methods mainly fall into two groups based on the diffusion and autoregressive paradigms. Diffusion-based approaches can be further categorized by their generation strategy: one-shot-for-all methods such as PartCrafter [18], PartPacker [31], OmniPart [40], BANG [43], FullPart [12] and X-Part [37], which jointly generate all parts in a single forward pass; one-shot-per-part models such as HoloPart [38] generates each part independently but conditioned on global context, and AutoPartGen [4] adopts a sequential generation scheme. In contrast, AR-based methods operate directly on mesh sequences: MeshCoder [7] leverages an LLM to produce Blender scripts that are executed to reconstruct parts, and Mesh Silksong [29] employs a manifold-aware mesh representation for autoregressive part synthesis. Except for adopting joint geometry-segmentation diffusion at the first stage, our part-level diffusion is also different in: (i) using latent-based conditions rather than explicit representations, which avoids geometry degradation, provides domain-aligned conditioning signals, and also injects informative part–whole relational cues; (ii) performing a dual-space latent generation procedure that achieves full-resolution part geometry of high quality and enables seamless and consistent part composition.
3 Method
Figure 1: The pipeline of UniPart. It includes a Geom-Seg VAE that encodes both whole geometry and part segmentation information into a unified representation, Geom-Seg VecSet. The image-guided part-level generation adopts a two-level pipeline, where a whole-level DiT first generates the whole geometry and segmented part latent, and a part-level DiT then accepts the input image and the whole-part latent as conditions for dual-space part latent generation. The final object mesh is composed of each full-resolution part mesh.
The task of part-level 3D shape generation aims to produce a decomposable 3D mesh $\mathcal{O}=\{\mathcal{M}_{i}\}_{i=1}^{N}$ from a prompt RGB image $I$, where each part of the object is represented by an individual mesh $\mathcal{M}$ and the number of parts $N$ is usually uncertain, depending on the complexity of the target object. The architecture of the proposed UniPart is illustrated in Fig. 1. It consists of (i) a geometry-segmentation variational autoencoder that represents both the geometry and segmentation information of each object mesh in the latent space; (ii) a global-level diffusion model that generates the global geometry attached with part segmentation information in the latent space; and (iii) a part-level diffusion model that synthesizes the fine-grained part geometry by conditioning on each part latent segment and the global geometry latent. Finally, all individual parts compose the target object mesh.
3.1 Geometry-Segmentation VAE
Preliminary: VecSet
[42] is a popular representation that encodes the geometry of a 3D mesh into the latent space. Specifically, dense points $P\in\mathbb{R}^{C\times 6}$ (each of $C$ points is represented by its position and normal vector on the mesh surface) are first sampled from the mesh $\mathcal{O}$. A fixed-length set of latent vectors $Z\in\mathbb{R}^{L\times d}$ is computed through a cross-attention layer between $P$ and its downsampled $L$ points (or a learnable query set) and several self-attention layers, encoded into $d$-dimension latent vectors. This process is conducted by the encoder $\mathcal{E}$ of a transformer-based VAE. The decoder $\mathcal{D}$ then decodes $Z$ into an implicit field (e.g. occupancy field or signed-distance field) for mesh reconstruction, as $\mathcal{D}(Z,q)\rightarrow[0,1]$, where $q$ is the set of query points in the target 3D space. The whole VAE is trained by minimizing the loss function formulated as
$\mathcal{L}_{\text{vecset}}=\mathcal{L}_{\text{recon}}+\lambda_{\text{kl}}\cdot\mathcal{L}_{\text{kl}},$ (1)
where $\mathcal{L}_{\text{recon}}$ is computed as $\|\mathcal{D}(Z,q)-f(q)\|_{2}^{2}$ with $f$ denoting the ground truth implicit field, $L_{\text{kl}}$ is the KL penalty for the divergence between $Z$ and Gaussian distribution, and $\lambda_{\text{kl}}>0$ is a weighting factor.
Geom-Seg VecSet
Classic VecSet representations encode rich geometric information into the latent space. To better support compositional generation, we further propose Geom-Seg Vecset that additionally encodes part segmentation information into the latent space and conducts latent segmenting in the decoding stage for part-wise representation. In doing so, we first segment object meshes into $N$ parts according to both geometric and semantic patterns. Then $C$ points with part segmentation label are sampled, denoted as $P\in\mathbb{R}^{C\times 7}$, as the input of encoding, derieving the encoded latent $Z=\mathcal{E}(P)\in\mathbb{R}^{L\times d}$. A decoder $\mathcal{D}_{\text{geom}}$ as in VecSet is adopted for geometry reconstruction from $Z$. We further develop a segmentation decoder $\mathcal{D}_{\text{seg}}$ and follow the promptable segmentation pipeline of SAM2 [27], which accepts $Z$ and latent indices $q$ as input and generates segmentation masks that the prompt belongs to. The Geom-Seg VecSet VAE is trained by minimizing the loss function formulated as
$\mathcal{L}_{\text{vecset}}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{seg}}+\lambda_{\text{kl}}\cdot\mathcal{L}_{\text{kl}},$ (2)
where $L_{\text{seg}}$ is the segmentation loss following SAM2 [27]. Considering geometric cues often delineate semantic boundaries, the geometric features and part segmentation patterns of objects are inherently interdependent. It critically facilitates coupled representation learning of geometry and segmentation information. In practice, we found that the Geom-Seg VecSet space can be smoothly learned by fine-tuning a pure geometric VecSet VAE, with the latent space and neural network kept in the original scale (equal in $L$ and $d$), which is sufficient to jointly represent the segmentation and geometry without loss of quality (see visualizations in Sec. 4.3).
3.2 Whole-level Latent Diffusion
With the learned geometry-segmentation latent space, we propose a two-stage pipeline for decomposable part-level generation, which consists of a whole-level diffusion transformer (DiT) [25] that generates the global geometry-segmentation latent and a part-level DiT to refine each part latent for complete part geometry of quality.
Geom-Seg Latent Diffusion
Our global-level DiT employs a rectified flow model [19], which uses a linear interpolation forward process $Z_{t}=(1-t)Z_{0}+t\epsilon$ to construct the noisy Geom-Seg VecSet latent $Z_{t}$ from the data sample $Z_{0}$ and noise $\epsilon\in\mathbb{R}^{L\times d}$ at a timestep $t\in[0,1]$. In the backward process, our DiT parameterized by $\theta$ learns to model the timestep-dependent velocity field, $v(Z_{t},t)=\delta_{t}Z_{t}$, representing the moving velocity from the noise to the data sample. $\theta$ is optimized on the conditional flow matching (CFM) objective [19]:
$\mathcal{L}_{\text{cfm}}(\theta)=\mathbb{E}_{t,Z_{0},\epsilon}\|v_{\theta}(Z_{t},t|I)-(\epsilon-Z_{0})\|_{2}^{2}.$ (3)
At the inference stage, we start from a randomly initialized noise and iteratively predict the velocity as $t$ minimized from 1 to 0, which progressively denoises $Z_{t}$ and derives the final global latent prediction $\hat{Z_{0}}$.
Latent Segmentation
[…]
3.3 Part-level Latent Diffusion
Given the generated whole-object latent $\hat{Z_{0}}$ and a set of segmented part latents $\{X_{i}\}_{i=1}^{N}$, we directly use them as condition signals to provide both global object geometry and local part information to synthesize high-quality part meshes.
Dual-Space Diffusion
Our part-level latent diffusion model adopts a dual-space generation architecture that jointly generates each part’s geometric latent from two spaces. For the $i$-th part, we denote its dual-space latent as $X_{i}^{*}\coloneqq(X_{i}^{\text{gcs}},X_{i}^{\text{ncs}})\in\mathbb{R}^{L\times 2d}$. Specifically, $X_{i}^{\text{gcs}}$ denotes the part latent in the global coordinate space, encoding the part’s relative scale, position, and compositional relationship to the whole object. In contrast, $X_{i}^{\text{ncs}}$ denotes the part latent in a normalized canonical space, capturing high-fidelity part geometry within a normalized $0,1]^{3}$ coordinate frame and sharing the same resolution as the whole-object latent (see illustration in Fig. 1). During dual-space diffusion training, a DiT parameterized by $\omega$ predicts $X_{i}^{*}$ for the $i$-th part, conditioned on the input image $I$, the part latent $X_{i}$ (padded to length $L$), and the whole-object latent $Z_{0}$ (replaced with $\hat{Z_{0}}$ at inference time). The prediction follows the same flow-matching formulation as in the whole-level latent diffusion, so we omit the detailed equations for brevity. To enable the transformer to distinguish tokens from the two geometric contexts, we inject a learnable space embedding $e^{s}\in\mathbb{R}^{1\times d}$, where $s\in\{\text{gcs},\text{ncs}\}$ indexes the coordinate space. The corresponding space embedding is added to $X_{i}^{\text{gcs}}$ and $X_{i}^{\text{ncs}}$ with broadcasting before feeding them into the transformer. Furthermore, we also adopt a global-local attention mechanism inspired by previous work [5, 37]:
$\text{Attn}_{\text{local}}=\text{Softmax}(\frac{\sigma_{q}(X_{i}^{s})^{\top}\sigma_{k}({X_{i}^{s}})}{\sqrt{h}})\sigma_{v}(X_{i}^{s}),$ (4)
$\text{Attn}_{\text{global}}=\text{Softmax}(\frac{\sigma_{q}(X_{i}^{*})^{\top}\sigma_{k}(X_{i}^{*})}{\sqrt{h}})\sigma_{v}(X_{i}^{*}),$ (5)
where $s\in{\text{gcs},\text{ncs}}$ indexes the coordinate space, $\sigma_{q/k/v}$ denotes the linear projections for query, key, and value, $h$ is the channel dimension of each attention head, and the diffusion timestep $t$ is omitted for simplicity. This global–local attention mechanism encourages joint attention to both whole-object and part-level geometry, leading to higher-quality part generation. In particular, $X_{i}^{\text{gcs}}$ shares the same coordinate system as the conditioning latents and can be viewed as a geometric completion of $X_{i}$, making it easier to predict. On the other hand, $X_{i}^{\text{gcs}}$ and $X_{i}^{\text{ncs}}$ are highly correlated, sharing rich part-level geometric information. These properties allow the coupled dual-space diffusion mechanism to learn high-quality geometry of each part in its own normalized canonical space more efficiently.
Part Decoding and Composition
At the inference stage, we iteratively denoise two random noise conditioned on $\hat{Z_{0}}$ and $X_{i}$ to get the denoised part latents in two spaces $X_{i}^{\text{gcs}}$ and $X_{i}^{\text{ncs}}$ for the $i$-th part. Using a shared pre-trained geometry decoder, we decode the latents into $\mathcal{M}_{i}^{\text{gcs}}$ and $\mathcal{M}_{i}^{\text{ncs}}$, respectively. In the final step, the scale and position of the $i$-th part mesh can be computed from $\mathcal{M}_{i}^{\text{gcs}}$, which is utilized to transform $\mathcal{M}_{i}^{\text{ncs}}$ of high geometric quality into the global coordinate space for part mesh composition.
4 Experiments
4.1 Implementation Details
Dataset Curation
We construct a curated dataset by integrating multiple public sources [9, 10, 6], yielding 302K objects (184K + 118K) with part-level segmentations. The 118K subset comprises exclusively watertight meshes. Part annotations are derived from mesh connectivity, with each object containing more than 1 part. To ensure topological consistency for SDF computation, all meshes are remeshed using a winding-number-based method [2]. Part labels are then propagated to the remeshed geometry by assigning each new face the label of its nearest original face.
Experiment Setting
We construct a test set of 100 objects with part-level annotations, each including rendered images and ground-truth part geometries. Since there is no reliable correspondence between generated parts and ground-truth parts, we concatenate all parts into a single unified mesh for evaluation to enable a holistic assessment of geometric fidelity. We report Chamfer Distance (CD) and F-Score at thresholds of 0.05 and 0.10 to evaluate fine- and coarse-level geometric alignment, respectively. All meshes are normalized to [1,1][-1,1], and the best score across rotations of $0^{\circ}$, $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$ is selected to ensure a pose-agnostic comparison. We implement five baseline methods for comparison: (i) PartPacker [31], (ii) PartCrafter [18], (iii) OmniPart [40], (iv) HoloPart [38], and (v) X-Part [37]. Among them, methods (i, ii, iii) accept single-image input for generation, the same as our approach. Methods (iv, v) require mesh input, so we generate the holistic object mesh using Huanyuan3D-2.1 [32], the same as ours, to enable fair comparison. Note that all baseline methods are implemented with the open-sourced version for evaluation.
Implementation Details
For the Geom-Seg VAE and global-level shape generation, we build upon Hunyuan3D-2.1 [32], using classifier-free guidance (CFG) with a dropout rate of 0.1 and the AdamW optimizer with a weight decay of 0.01. For part-level generation, we initialize the weights from Hunyuan3D-2.1 and fine-tune with the same AdamW settings. All training is conducted on 8 NVIDIA A800 GPUs (80GB) with a batch size of 32. More implementation details are included in the supplementary materials.
4.2 Experiment Results
Geom-Seg VAE Reconstruction
We first present the reconstruction performance of our proposed Geom-Seg VAE, as shown in Fig. 2. Both the reconstructed geometry and the inferred part-level latent segmentation demonstrate high fidelity to the input, preserving overall shape and part details while maintaining coherent part assignments across the object. We also provide results from Hunyuan3D-2.1 [32] as a reference for the reconstructed geometry. These reference reconstructions offer a baseline for comparison, illustrating that our Geom-Seg VAE achieves accurate part-level segmentation while maintaining high-quality geometry. As an intermediate product, the whole-object geometry and part latent segmentation actually provide a solid support for the part-level diffusion at the second stage.
Figure 2: Reconstruction results of our Geom-Seg VAE. (a) Input mesh with part segmentation masks; (b) Reconstructed geometry; (c) Reconstructed part latent segmentation visualization. (d) Reconstructed geometry of Hunyuan3D-2.1 [32] for reference. Better zoom in for details.
Whole-level Generation Results
Fig. 3 visualizes some sampled outputs by the first-stage whole-level diffusion model to demonstrate its effectiveness. Accepting various types of object images as input, our whole-level DiT generates high-quality global geometry together with consistent part-level latent segmentations, indicating reliable synthesis of both object shape and part segmentation.
Figure 3: Generated results of our whole-level DiT. (a) Input image; (b) Generated whole-object geometry; (c) Generated part latent segmentation. Better zoom in for details.
Part-level Generation Results
We present both quantitative and qualitative evaluations of the final part-level generation produced by UniPart, and compare against baseline methods, including HoloPart [38], OmniPart [40], PartCrafter [18], PartPacker [31], and XPart [37]. Qualitative results are provided in Tab. 2. Our UniPart outperforms all other competitors consistently on all metrics (i.e., CD and F-Score with two distinct thresholds), which shows our method produces final objects with higher geometric quality and better alignment to the ground truth. We further provide qualitative results in Fig. 4, which illustrates that our UniPart can generate more reasonable part segmentations and higher-quality part geometries. Some other part-level generations are visualized in Fig. 5. More visual results are shown in the teaser figure and supplementary materials.
Figure 4: Qualitative results of part-level 3D generation for a given image. We visualize the “exploded” parts in the first line of each mesh pair for better visualization of part generation. Our UniPart can produce more reasonable part segmentations and higher-quality part geometries.
Table 1: Quantitative comparison on the geometry quality of part-level generation results. Best results are marked in bold font. Our UniPart significantly outperforms other methods on both CD and F-score metrics.
Method        CD $\downarrow$     F@0.05 $\uparrow$     F@0.10 $\uparrow$
HoloPart [38] 0.1492 0.5208 0.7450
OmniPart [40] 0.1453 0.5273 0.7656
PartCrafter [18] 0.1778 0.4749 0.7120
PartPacker [31] 0.1654 0.4715 0.7226
XPart [37] 0.1533 0.5242 0.7523
UniPart (Ours) 0.1311 0.5565 0.8052
Figure 5: More generation results by our UniPart. (a) Input images; (b) The whole object geometry output by our whole-level DiT; (c) The composed part meshes output by our part-level DiT; (d) The part latent segmentation visualization produced by our whole-level DiT.
Comparison on 3D Segmentation
Our UniPart actually conducts part segmentation during generation, so the segmentation performance can not be directly evaluated on existing 3D part segmentation benchmarks. Therefore, we specifically construct an evaluation set with 50 object meshes generated by our whole-level DiT (decode the global-geometry latent with Geom. decoder). A group of ten human annotators is invited to manually annotate the part segmentation on these meshes, also given the input image as a reference. The collected annotations are further double-checked and finely corrected by annotators in the second round to guarantee high quality. As a result, we obtain 536 high-quality part annotations from the 50 object meshes. Based on them, we evaluate the part segmentation results of UniPart with existing SOTA baseline methods, including SAMesh [30], PartField [20], and P3-SAM [23]. For a direct comparison with ours, which only contains 4,096 latent vectors as part segmentation, we uniformly sample 4,096 points from each segmented point cloud or mesh surface of other methods to compute the numerical results. For quantitative results, we adopt mean IoU (mIoU) as the evaluation metric, with results shown in Tab. 2. We found that when the geometry pattern on the mesh surface is not significant enough to support feed-forward-style part segmentation, our diffusion-based generative segmentation can still achieve robust performance, thanks to the part knowledge gained in the generative geometry learning process.
Table 2: Quantitative comparison on the geometry quality of part-level generation results. Best results are marked in bold font.
Method SAMesh [30] PartField [20] P3-SAM [23] UniPart (Ours)
mIoU $\uparrow$ 0.3608 0.4167 0.7046 0.7222
4.3 Ablation Studies
Figure 6: Visual results of ablation studies for the design of UniPart, i.e., without using normalized canonical space (NCS) generation, local attention, and space embedding injection.
Figure 7: Visual results of ablation studies for the design of UniPart, including using global geometry condition and conditioning the part-level DiT in the form of latent.
Comparison between the geometric reconstruction of Geom-Seg VecSet VAE and original VecSet VAE (Hunyuan3D-2.1 VAE) has been shown in Fig. 2, which proves that adding segmentation information into the latent space does not result in a drop in reconstruction quality. The learning of the whole-level DiT follows a typical latent diffusion model as in previous methods. Thus, we focus on the ablation study on the design of part-level DiT. There are three key designs: (i) dual-space generation for part meshes in the normalized canonical space (NCS), rather than directly using the global coordination generation; (ii) local-global attention mechanism in the dual-space diffusion; (iii) space embedding injection for distinguishing tokens from different spaces within the attention mechanism to provide clearer guidance. We visualize the impact of each component using two examples shown in Fig. 6. The results validate that generation in the normalized canonical space effectively enhances part geometric fidelity, especially for small parts with fine details. Local attention promotes part cohesion, yielding more harmonious spatial arrangements and improved global geometric consistency. Without space embedding injection, the model frequently confuses the target space for each part, resulting in catastrophic failures during assembly.
In our Part-level DiT design, we feed global geometry information as a condition to enable whole-aware part generation. A visual comparison of not using the global geometry condition is shown in Fig. 7. Compared with the full version, removing the global geometry condition influences the seamlessness of the composition. Besides, we also provide a comparison between using latent-form conditions and using classic point-form conditions. In previous methods like X-Part [37] and HoloPart [38], they sample dense points from the segmented object mesh and further train feature extraction layers for geometry encoding, which actually introduces (i) information loss in sampling and encoding; (ii) a domain gap with latent vectors to denoise; and (iii) extra training burden. As shown in Fig. 7, using point-form conditions results in geometry degradation at a fine-grained level, especially for some small-scale parts.
5 Conclusion
We introduced a unified framework for part-level 3D generation that leverages a geometry–segmentation latent space and a dual-stage latent diffusion process. By constructing Geom-Seg VecSet, a joint latent representation that encodes both whole-object geometry and part-level structure, our method enables part-aware understanding to emerge naturally from whole-level generative learning without relying on external segmenters. Built on this representation, our diffusion model UniPart performs joint geometry generation and part-level latent segmentation, followed by a dual-space diffusion that predicts part latents in both global and canonical coordinate spaces for high-fidelity reconstruction and consistent composition. Extensive experiments demonstrate that our approach outperforms existing methods in segmentation controllability, geometric fidelity, and overall part-level generation quality. We believe that the proposed unified latent space and dual-space diffusion framework pave the way for future research in structured 3D generation, editable 3D content creation, and part-aware generative modeling.
Limitations and Future Work
First, our UniPart can only handle decomposable object generation, lacking validation on the ability of scene generation. This will be our future work to explore. Second, our Geom-Seg VecSet is constructed based on point-based latent representation, VecSet, and we have not explored implementing the paradigm based on voxel-based latent representation, like SLAT, which is also one of our future works.